OpenStudy (anonymous):

Determine the perimeter of quadrilateral KLMN with vertices K(-4, 2), L(2, 2), M(8, -3), and N(-10, -6).

OpenStudy (shubhamsrg):

i dont think theres any shortcut here..you'll have to use distance formulla here.. 4 times here it is,,distance between (x1,y1) and (x2,y2) is sqrt[(x2-x1)^2 + (y2-y1)^2] hope this helps..

OpenStudy (anonymous):

I did.. but I am getting...48.1

OpenStudy (amistre64):

can you show your work and we might be able to see where, or if, its got an error in it

OpenStudy (anonymous):

KL=6, LM=7.8, MN=18.2, NK=16.12

OpenStudy (amistre64):

L(2, 2), M(8, -3) -2-2 -2-2 --------------- 0,0 6,-5; sqrt(36+25) sqrt(61) , good-ish

OpenStudy (amistre64):

how accurate does the solution have to be?

OpenStudy (anonymous):

round it

OpenStudy (anonymous):

it doesnt have to be too acruarate

OpenStudy (amistre64):

M(8, -3), N(-10, -6) 10 6 10 6 ------------------- 18, 3 0,0 sqrt(324+9) sqrt(333) = 18.248 , that one checks out too

OpenStudy (amistre64):

N = -10,-6 -K = 4, -2 ------------ -6,-8; sqrt(36+64) sqrt(100) = 10 , that one looks off

OpenStudy (amistre64):

subtract 6.12 from your answer and youll should be closer

OpenStudy (anonymous):

answer is 42

OpenStudy (amistre64):

yep, 48.1 - 6.1 = 42

OpenStudy (anonymous):

THANK U

OpenStudy (amistre64):

youre welcome